# Regional and partial observability and control of waves

### v1

@@ 3.1. Local geodesic coordinates
Near a point $m_{0}$ of the boundary $\partial\Omega$, taking advantage of the regularity of $\Omega$ which is of class $\mathcal{C}^{\infty}$, we can define a system of geodesic coordinates $x=(x_{1},x_{2},....,x_{n})\mapsto y=(y_{1},y_{2},....,y_{n})$ such that we have locally
@@
Here, the matrix $(b_{ij}(y))_{ij}$ is of class $\mathcal{C}^{\infty}$, symmetric, uniformly definite positive on a neighborhood of $m_{0}$, $M_{0}(y)$ is a real valued function of class $\mathcal{C}^{\infty}$, and $M_{1}(y,\partial_{y^{\prime}})$ is a tangential differential operator of order $1$ with $\mathcal{C}^{\infty}$ coefficients.
@@ 3.2. Generalized bicharacteristic rays
(3.4) #$\#\Big{\{}\pi^{-1}(\rho)\cap\text{Char}(P_{A})\Big{\}}=\left\{\begin{array}[]{c}0\quad if\quad\rho\in\mathcal{E}\\ 1\quad if\quad\rho\in\mathcal{G}\\ 2\quad if\quad\rho\in\mathcal{H}.\end{array}\right.$
@@
=$\mathcal{E}=\{r_{0}<0\},\quad\quad\mathcal{H}=\{r_{0}>0\},\quad\quad\mathcal{G}=\{r_{0}=0\}.$
@@
Let $\rho$ be a point of $\0T^{\ast}\partial\mathcal{L}\backslash 0$. We say that
@@
    This means that the free bicharacteristic ray $\gamma$ issued from $\rho$ belongs to $T^{\ast}\mathcal{L}$ in a neighborhood of 00, except at $s=0$, i.e., there exists $\varepsilon>0$ such that $(\exp sH_{p_{A}})\widetilde{\rho}\in T^{\ast}\mathcal{L}$ for $0<|s|<\varepsilon$, with $\widetilde{\rho}=\pi^{-1}(\rho)$.
@@
    $\rho$ is nondiffractive if a) $\rho\in\mathcal{H}$ or b) $\rho\in\mathcal{G}$ and the free bicharacteristic ray $(\exp sH_{p_{A}})\widetilde{\rho}$ passes over the complement of $\mathcal{L}$ for arbitrarily small values of $s$
@@
    In the interior, i.e in $T^{*}\mathcal{L}$, a generalized bicharacteristic is simply a classical bicharacteristic ray of the wave operator whose projection on the space is a geodesic of $\Omega$ equipped with the metric $(a^{ij})=(a_{ij})^{-1}$
@@ 3.3. Sets of interest
(3.6) $\mathscr{R}(\omega_{T})$ =$=\{\rho=(t,x,\tau,\xi)\in T^{*}_{b}\mathcal{L}\backslash 0,\text{s. t. }\rho\notin\Sigma_{A}\text{ or }\gamma_{\rho}({\mathbb{R}})\cap T^{*}_{b}\omega_{T}\neq\emptyset\},$
@@
(3.7) $\mathscr{R}_{0}(\omega_{T})=\Big{\{}(x,\xi)\in T_{b}^{*}\Omega\backslash 0,\\ \text{ s. t. any $\gamma_{\rho}$}\text{ emanating from $(x,\xi)$ at $t=0$ satisfies }\gamma_{\rho}\cap T^{*}_{b}(\omega_{T})\neq\emptyset\Big{\}}.$
@@
    b) If $|\tau|>|\xi^{\prime}|_{x}$, we are dealing with a hyperbolic point. $\gamma$ is then one of the two hyperbolic fibers of $P_{A}$ at $\rho$. According to the hamiltonian equations, on sees that the bicharacteristic curve $\gamma$ corresponds to the integral curve of the (free) hamiltonian field $H_{p}$ issued from the point $\rho_{-}=\big{(}0,x;\tau<-|\xi^{\prime}|_{x},\xi^{\prime},\xi_{n}=+\sqrt{\tau^{2}-|\xi^{\prime}|_{x}^{2}}\,\big{)}$ or $\rho_{+}=\big{(}0,x;\tau>|\xi^{\prime}|_{x},\xi^{\prime},\xi_{n}=+\sqrt{\tau^{2}-|\xi^{\prime}|_{x}^{2}}\,\big{)}$.
@@
The geometric condition of Theorem 1.1 can in fact be simply stated as $\overline{\mathscr{O}}\subset\mathcal{O}(\omega_{T})$. In other words, Theorem 1.1 applies for any open set $\mathscr{O}$ strictly included in $\mathcal{O}(\omega_{T})$ with the observation set $\omega_{T}$.
@@ 3.4. Pseudo-differential operators
Following [11] ,we define the set $\mathcal{A}$ of pseudo-differential operators on ×${\mathbb{R}}\times{\mathbb{R}}^{n}$ of the form $Q=Q_{i}+Q_{\partial}$ where $Q_{i}$ is a classical pseudo-differential operator, compactly supported in $\mathcal{L}$ and $Q_{\partial}$ is a classical pseudo-differential operator tangential to the boundary $\partial\mathcal{L}$, compactly supported near $\partial\mathcal{L}$. More precisely, $Q_{i}=\varphi Q_{i}\varphi$ for some $\varphi\in\mathcal{C}^{\infty}_{0}(\mathcal{L})$, and $Q_{\partial}=\psi Q_{\partial}\psi$ for some $\psi\in\mathcal{C}_{0}^{\infty}(U_{\partial\mathcal{L}})$, where $U_{\partial\mathcal{L}}$ is a small neighborhood of $\partial\mathcal{L}$ in ×${\mathbb{R}}\times{\mathbb{R}}^{n}$. For $s\in{\mathbb{R}}$, $\mathcal{A}^{s}$ denotes the set of elements of order $s$ of $\mathcal{A}$.
In a similar way, we also define the set $\mathcal{B}$ of pseudo-differential operators on ${\mathbb{R}}^{n}$, i. e. on the space variable, of the form $\psi=\psi_{i}+\psi_{\partial}$ where $\psi_{i}$ is a classical pseudo-differential operator, compactly supported in $\Omega$, and $\psi_{\partial}$ is a classical pseudo-differential operator tangential to the boundary $\partial\Omega$, compactly supported near $\partial\Omega$. Similarly as above, for $s\in{\mathbb{R}}$, $\mathcal{B}^{s}$ denotes the set of the elements of $\mathcal{B}$ of order $s$.
@@ 3.5. Wave front sets and propagation results
For a distribution $u$ defined on the cylinder =×$\mathcal{L}=\mathbb{R}\times\Omega$, we define the $H^{s}$-wave front set up to the boundary, denoted $WF^{s}_{b}(u)$, as a subset of the compressed cotangent bundle in the sense of Melrose–Sjöstrand, $T^{*}_{b}\mathcal{L}=T^{\ast}\mathcal{L}\cup T^{\ast}\partial\mathcal{L}$. This set coincides with the classical wave front set $WF^{s}(u)$ in the interior of $\mathcal{L}$, i.e., in $T^{*}\mathcal{L}$, and extends the notion to describe the $H^{s}$-microlocal regularity of $u$ up to the boundary. We follow here the definition of Chazarain (see [3]) that is used by Melrose and Sjöstrand (see [15]). In addition, for solutions of $P_{A}u\in\mathcal{C}^{\infty}$, it agrees with the intrinsic notion of Melrose (see Hörmander [6, Cor.18.3.33]), which does not depend on $P_{A}$. In the following, we use the spaces of pseudodifferential operators $\mathcal{A}^{0}$ and $\mathcal{B}^{0}$, introduced in Section 3.4.
Consider $s\in{\mathbb{R}}$ and $u\in\mathcal{D}^{\prime}(\mathcal{L})$ solution to $P_{A}u=0$ in $\mathcal{L}$. Also, for $q\in{\mathbb{R}}^{n+1}$ and $r>0$, denote by $B_{r}(q)$ the Euclidean ball of center $q$ and radius $r$.
@@
  • If $\rho=(q,\eta)\in T^{\ast}\mathcal{L}$, i.e $q$ is an interior point, there exists a pseudodifferential operator $Q=Q_{i}\in\mathcal{A}^{0}$, elliptic at $\rho$, such that $Qu\in H^{s}(B_{r}(q))$ for some $r>0$ , $B_{r}(q)\subset\mathcal{L}$
  • If $\rho=(q,\eta)\in T^{\ast}\partial\mathcal{L}$, i.e $q$ is a boundary point, there exists a tangential pseudodifferential operator $Q=Q_{\partial}\in\mathcal{A}^{0}$, elliptic at $\rho$, such that $Qu\in H^{s}(B_{r}(q)\cap\mathcal{L})$ for some $r>0$.
@@
We will deduce that $u$ actually belongs to $H^{1}((0,T)\times\Omega)$. Indeed, let us consider $\rho_{0}\in T^{*}_{b}((0,T)\times\Omega)$. If $\rho_{0}$ is an elliptic point (independently if it is an interior or a boundary point), it is classical that $\rho_{0}\notin WF_{b}^{2}u$, i.e., $u$ is in $H^{2}$ microlocally near $\rho_{0}$. Here, a special care must be taken at the boundary, see Hörmander [6, Theorem 20.1.14] .
@@
Accordingly, the solution $u$ of the wave system (1.1) lies in $H^{1}((0,T)\times\Omega)$. Thus, by conservation of energy in time, we also conclude that the initial data has finite energy, i.e.,$(u_{0},u_{1})$ belongs to $H^{1}_{0}(\Omega)\times L^{2}(\Omega)$. $\square$
@@
The proof of the propagation of regularity from the boundary stated in Lemma 2.5 follows the same strategy as the one of Lemma 2.1. The only difference is that one has to use the propagation of the $H^{1}$-wave front set from an observation on the boundary, namely out of the information that $\partial_{n}u_{|\partial\Omega}\in L^{2}(\Gamma_{T})$. This can be done on nondiffractive points. Indeed, under the assumption $\partial_{n}u_{|\partial\Omega}\in L^{2}(\Gamma_{T})$, by the lifting lemma in [1, Theorem 2.2], we deduce that every nondiffractive point $\rho_{0}\in T^{*}(\Gamma_{T})$ satisfies $\rho_{0}\notin WF_{b}^{1}(u)$. In other words, the solution $u$ is $H^{1}$ microlocally near this point, up to the boundary. This suffices to conclude the proof of Lemma 2.5 in view of the imposed microlocal geometric condition on the pair $(\Gamma,\mathscr{O})$.
@@ 3.7. Further technical results
Let $x_{0}\in\mathcal{O}(\omega_{T})$. Since $\mathcal{O}(\omega_{T})$ is open in $\overline{\Omega}$, by continuity of the flow of Melrose-Sjöstrand, there exists $\varepsilon>0$ and $\varphi\in C^{\infty}_{c}((-\varepsilon,\varepsilon)\times\overline{\Omega})$ such that $\varphi\equiv 1$ in a neighborhood of $(0,x_{0})$, such that all rays emanating from the support of $\varphi$ meet $\omega_{T}$.
@@
Our first goal is to prove that for any $\psi(x,D_{x})\in\mathcal{B}_{0}$ supported in $\mathscr{R}_{0}(\omega_{T})$, we have $\psi(x,D_{x})u\in H^{1}((0,\varepsilon)\times\Omega)$. And for this end, we will prove that any point $\rho_{0}=(t_{0},x_{0};\tau_{0},\xi_{0})\in T^{*}_{b}\mathcal{L},t_{0}\in]0,\varepsilon[$, satisfies $\rho_{0}\notin WF^{1}_{b}(\psi(x,D_{x})u)$.
Case 1: $\rho_{0}\in T^{*}\mathcal{L}$, i.e., it is an interior point and $x_{0}\in\Omega$. Take $\varphi=\varphi(t,x)\in\mathcal{C}_{0}^{\infty}((0,\varepsilon)\times\Omega)$, supported near $(T,x_{0})$. In the operators space $\mathcal{A}_{0}$ consider the local identity partition
@@
Clearly, $\operatorname{Supp}(q_{2})$ is contained in the elliptic set of $T^{*}\mathcal{L}$, so $q_{2}(t,x;D_{t},D_{x})u\in H^{1}(\mathcal{L}_{\varepsilon})$, and $\psi(x,D_{x})q_{2}u\in H^{1}(\mathcal{L}_{\varepsilon})$
Let us now examine the first term $q_{1}(t,x;D_{t},D_{x})u$. Here we notice that the composition $\psi(x,D_{x})q_{1}(t,x;D_{t},D_{x})$ provides a well defined global pseudodifferential operator, according to [6, Th. 18.1.35]. In addition, if $\rho=(t,x;\tau,\xi)\in\operatorname{Supp}(\sigma(\psi q_{1}))$, where $\sigma$ denotes the symbol of the operator $\psi q_{1}$, $\rho$ is either an elliptic or hyperbolic point lying in the set $E_{\varepsilon}$, which does not intersect $WF^{1}u$. Therefore, $\psi(x,D_{x})q_{1}(t,x;D_{t},D_{x})u\in H^{1}(\mathcal{L}_{\varepsilon})$.
@@
Here we shall work in a system of local geodesic coordinates $(t,x^{\prime},x_{n};\tau,\xi^{\prime},\xi_{n})$ with $\partial\Omega=\{x_{n}=0\}$ and $\Omega=\{x_{n}>0\}$, see Section 3.1. Hence we will set $\rho_{0}=(T,x^{\prime}_{0},\tau_{0},\xi^{\prime}_{0})$. Recall that the operators of $\mathcal{A}_{0}$ (resp. of $\mathcal{B}_{0}$) take the form $q(x_{n},t,x^{\prime},D_{t},D_{x^{\prime}})$ (resp. $\psi(x_{n},x^{\prime},D_{x^{\prime}})$).
As in Case 1 above, we consider $\varphi=\varphi(t,x)\in\mathcal{C}_{0}^{\infty}((0,\varepsilon)\times{\mathbb{R}}^{n})$, supported near $(T,x_{0})$, and a local partition of the identity with tangential pseudodifferential operators, of the form
@@
With notations of Section 3.2, $\operatorname{Supp}(q_{2})\subset\mathcal{H}$, the hyperbolic subset of $T^{*}\partial\mathcal{L}$. Therefore since $\operatorname{Supp}(\psi)\subset\mathscr{R}_{0}(\omega_{T})$, we get $q_{2}(x_{n},t,x^{\prime};D_{t},D_{x^{\prime}})\psi(x_{n},x^{\prime},D_{x^{\prime}})u\in H^{1}(\mathcal{L}_{T})$
Finally, as in Case 1 , we notice that the composition $q_{1}(x_{n},t,x^{\prime};D_{t},D_{x^{\prime}})\psi(x_{n},x^{\prime},D_{x^{\prime}})$ provides a well defined global tangential pseudodifferential operator, see [6, Th. 18.1.35], whose support is contained in $\mathcal{E}\cup\mathscr{R}(\omega_{T})$ And this yields $q_{1}(x_{n},t,x^{\prime};D_{t},D_{x^{\prime}})\psi(x_{n},x^{\prime},D_{x^{\prime}})u\in H^{1}(\mathcal{L}_{T})$ according to item (1).
Let us now examine the regularity of the trace $(u_{0},u_{1})$. For this, consider a function $h(t)\in\mathcal{C}^{\infty}_{0}({\mathbb{R}})$, $h(t)=1$ for $|t|\leqslant\varepsilon/2$ and $h(t)=0$ for $|t|\geqslant 3/4\varepsilon$. For $u(t,x)$ solution to (1.1), the function $v=h(t)\psi(x,D_{x})u$ satisfies the wave system
@@
The right hand side of this equation lies in $L^{2}((0,\varepsilon)\times\Omega)$ according to the argument above. Therefore $(v(0,\cdot),\partial_{t}v(0,\cdot))=\psi(x,D_{x})(u_{0},u_{1})$ belongs to $H^{1}_{0}(\Omega)\times L^{2}(\Omega)$. This ends the proof of item (2) in Lemma 3.7, since $\psi(x,D)$ is any operator in $\mathcal{B}_{0}$ supported in $\mathscr{R}_{0}(\omega_{T})$. ∎

### v2

@@ 3.1. Local geodesic coordinates
Near a point $m_{0}$ of the boundary $\partial\Omega$, taking advantage of the regularity of $\Omega$ which is of class $\mathscr{C}^{\infty}$, we can define a system of geodesic coordinates $x=(x_{1},x_{2},....,x_{n})\mapsto y=(y_{1},y_{2},....,y_{n})$ such that we have locally
@@
Here, the matrix $(b_{ij}(y))_{ij}$ is of class $\mathscr{C}^{\infty}$, symmetric, uniformly definite positive on a neighborhood of $m_{0}$, $M_{0}(y)$ is a real valued function of class $\mathscr{C}^{\infty}$, and $M_{1}(y,\partial_{y^{\prime}})$ is a tangential differential operator of order $1$ with $\mathscr{C}^{\infty}$ coefficients.
@@ 3.2. Generalized bicharacteristic rays
(3.4) $\Big\{\pi^{-1}(\rho)\cap\text{Char}(P_{A})\Big\}=\left\{\begin{array}[]{c}0\quad if\quad\rho\in\mathcal{E}\\ 1\quad if\quad\rho\in\mathcal{G}\\ 2\quad if\quad\rho\in\mathcal{H}.\end{array}\right$.
@@
$\mathcal{E}=\{r_{0}<0\},\quad\quad\mathcal{H}=\{r_{0}>0\},\quad\quad\mathcal{G}=\{r_{0}=0\}.$
@@
Let $\rho$ be a point of $T^{\ast}\partial\mathcal{L}\backslash 0$. We say that
@@
    This means that the free bicharacteristic ray $\gamma$ issued from $\rho$ belongs to $T^{\ast}\mathcal{L}$ in a neighborhood of 0, except at $s=0$, i.e., there exists $\varepsilon>0$ such that $(\exp sH_{p_{A}})\widetilde{\rho}\in T^{\ast}\mathcal{L}$ for $0<|s|<\varepsilon$, with $\widetilde{\rho}=\pi^{-1}(\rho)$.
@@
    $\rho$ is nondiffractive if a) $\rho\in\mathcal{H}$ or b) $\rho\in\mathcal{G}$ and the free bicharacteristic ray $(\exp sH_{p_{A}})\widetilde{\rho}$ passes over the complement of $\mathcal{L}$ for arbitrarily small values of $s$ .
@@
    In the interior, i.e., in $T^{*}\mathcal{L}$, a generalized bicharacteristic is simply a classical bicharacteristic ray of the wave operator whose projection on the space variable is a geodesic of $\Omega$ equipped with the metric $(a^{ij})=(a_{ij})^{-1}$ .
@@ 3.3. Sets of interest
(3.6) $\mathscr{R}(\omega_{T})$ $=\{\rho=(t,x,\tau,\xi)\in T^{*}_{b}\mathcal{L}\backslash 0,\text{ s. t. }\rho\notin\Sigma_{A}\text{ or }\gamma_{\rho}({\mathbb{R}})\cap T^{*}_{b}\omega_{T}\neq\emptyset\},$
@@
[…]
@@
    b) If $|\tau|>|\xi^{\prime}|_{x}$, we are dealing with a hyperbolic point. $\gamma$ is then one of the two hyperbolic fibers of $P_{A}$ at $\rho$. According to the hamiltonian equations, on sees that the bicharacteristic curve $\gamma$ corresponds to the integral curve of the (free) hamiltonian field $H_{p}$ issued from the point $\rho_{-}=\big(0,x;-|\tau|,\xi^{\prime},\xi_{n}=+\sqrt{\tau^{2}-|\xi^{\prime}|_{x}^{2}}\,\big)$ or $\rho_{+}=\big(0,x;+|\tau|,\xi^{\prime},\xi_{n}=+\sqrt{\tau^{2}-|\xi^{\prime}|_{x}^{2}}\,\big)$.
@@
The geometric condition of Theorem 1.1 can in fact be simply stated as $\overline{\mathscr{O}}\subset\mathcal{O}(\omega_{T})$. In other words, Theorem 1.1 applies for any open set $\mathscr{O}$ strictly included in $\mathcal{O}(\omega_{T})$.
@@ 3.4. Pseudo-differential operators
Following [12] ,we define the set $\mathcal{A}$ of pseudo-differential operators on ${\mathbb{R}}\times{\mathbb{R}}^{n}$ of the form $Q=Q_{i}+Q_{\partial}$ where $Q_{i}$ is a classical pseudo-differential operator, compactly supported in $\mathcal{L}$ and $Q_{\partial}$ is a classical pseudo-differential operator tangential to the boundary $\partial\mathcal{L}$, compactly supported near $\partial\mathcal{L}$. More precisely, $Q_{i}=\varphi Q_{i}\varphi$ for some $\varphi\in\mathscr{C}^{\infty}_{0}(\mathcal{L})$, and $Q_{\partial}=\psi Q_{\partial}\psi$ for some $\psi\in\mathscr{C}_{0}^{\infty}(U_{\partial\mathcal{L}})$, where $U_{\partial\mathcal{L}}$ is a small neighborhood of $\partial\mathcal{L}$ in ${\mathbb{R}}\times{\mathbb{R}}^{n}$. For $s\in{\mathbb{R}}$, $\mathcal{A}^{s}$ denotes the set of elements of order $s$ of $\mathcal{A}$.
In a similar way, we also define the set $\mathcal{B}$ of pseudo-differential operators on ${\mathbb{R}}^{n}$, i.e., on the space variable, of the form $\psi=\psi_{i}+\psi_{\partial}$ where $\psi_{i}$ is a classical pseudo-differential operator, compactly supported in $\Omega$, and $\psi_{\partial}$ is a classical pseudo-differential operator tangential to the boundary $\partial\Omega$, compactly supported near $\partial\Omega$. Similarly as above, for $\in{\mathbb{R}}, s\mathcal{B}^{s}$ denotes the set of the elements of $\mathcal{B}$ of order $s$.
@@ 3.5. Wave front sets and propagation results
For a distribution $u$ defined on the cylinder $\mathcal{L}=\mathbb{R}\times\Omega$, we define the $H^{s}$-wave front set up to the boundary, denoted $WF^{s}_{b}(u)$, as a subset of the compressed cotangent bundle in the sense of Melrose–Sjöstrand, $T^{*}_{b}\mathcal{L}=T^{\ast}\mathcal{L}\cup T^{\ast}\partial\mathcal{L}$. This set coincides with the classical wave front set $WF^{s}(u)$ in the interior of $\mathcal{L}$, i.e., in $T^{*}\mathcal{L}$, and extends the notion to describe the $H^{s}$-microlocal regularity of $u$ up to the boundary. We follow here the definition of Chazarain (see [4]) that is used by Melrose and Sjöstrand (see [16]). In addition, for solutions of $P_{A}u\in\mathscr{C}^{\infty}$, it agrees with the intrinsic notion of Melrose (see Hörmander [7, Cor.18.3.33]), which does not depend on $P_{A}$. In the following, we use the spaces of pseudodifferential operators $\mathcal{A}^{0}$ and $\mathcal{B}^{0}$, introduced in Section 3.4.
Consider $s\in{\mathbb{R}}$ and $u\in\mathscr{D}^{\prime}(\mathcal{L})$ (later, $u$ will be a solution of $P_{A}u=0$ in $\mathcal{L}$). Also, for $q\in{\mathbb{R}}^{n+1}$ and $r>0$, denote by $B_{r}(q)$ the Euclidean ball of center $q$ and radius $r$.
@@
  • If $\rho=(q,\eta)\in T^{\ast}\mathcal{L}$, i.e., $q$ is an interior point, there exists a pseudodifferential operator $Q=Q_{i}\in\mathcal{A}^{0}$, elliptic at $\rho$, such that $Qu\in H^{s}(B_{r}(q))$ for some $r>0$ , $B_{r}(q)\subset\mathcal{L}$ .
  • If $\rho=(q,\eta)\in T^{\ast}\partial\mathcal{L}$, i.e., $q$ is a boundary point, there exists a tangential pseudodifferential operator $Q=Q_{\partial}\in\mathcal{A}^{0}$, elliptic at $\rho$, such that $Qu\in H^{s}(B_{r}(q)\cap\mathcal{L})$ for some $r>0$.
@@
We will deduce that $u$ actually belongs to $H^{1}((0,T)\times\Omega)$. Indeed, let us consider $\rho_{0}\in T^{*}_{b}((0,T)\times\Omega)$. If $\rho_{0}$ is an elliptic point (independently if it is an interior or a boundary point), it is classical that $\rho_{0}\notin WF_{b}^{2}u$, i.e., $u$ is in $H^{2}$ microlocally near $\rho_{0}$. Here, a special care must be taken at the boundary, see Hörmander [7, Theorem 20.1.14] .
@@
[…]
@@
The proof of the propagation of regularity from the boundary stated in Lemma 2.5 follows the same strategy as the one of Lemma 2.1. The only difference is that one has to use the propagation of the $H^{1}$-wave front set from an observation on the boundary, namely out of the information that $\partial_{n}u_{|\partial\Omega}\in L^{2}(\Gamma_{T})$. This can be done on nondiffractive points. Indeed, under the assumption $\partial_{n}u_{|\partial\Omega}\in L^{2}(\Gamma_{T})$, by the lifting lemma in [2, Theorem 2.2], we deduce that every nondiffractive point $\rho_{0}\in T^{*}(\Gamma_{T})$ satisfies $\rho_{0}\notin WF_{b}^{1}(u)$. In other words, the solution $u$ is $H^{1}$ microlocally near this point, up to the boundary. This suffices to conclude the proof of Lemma 2.5 in view of the imposed microlocal geometric condition on the pair $(\Gamma,\mathscr{O})$.
@@ 3.7. Further technical results
Let $x_{0}\in\mathcal{O}(\omega_{T})$. Since $\mathcal{O}(\omega_{T})$ is open in $\overline{\Omega}$, by continuity of the flow of Melrose-Sjöstrand, there exists $\varepsilon>0$ and $\varphi\in\mathscr{C}^{\infty}_{c}((-\varepsilon,\varepsilon)\times\overline{\Omega})$ such that $\varphi\equiv 1$ in a neighborhood of $(0,x_{0})$, such that all rays emanating from the support of $\varphi$ meet $\omega_{T}$.
@@
Our first goal is to prove that for any $\psi(x,D_{x})\in\mathcal{B}^{0}$ supported in $\mathscr{R}_{0}(\omega_{T})$, we have $\psi(x,D_{x})u\in H^{1}((0,\varepsilon)\times\Omega)$. And for this end, we will prove that any point $\rho_{0}=(t_{0},x_{0};\tau_{0},\xi_{0})\in T^{*}_{b}\mathcal{L},t_{0}\in(0,\varepsilon)$, satisfies $\rho_{0}\notin WF^{1}_{b}(\psi(x,D_{x})u)$.
Case 1: $\rho_{0}\in T^{*}\mathcal{L}$, i.e., it is an interior point and $x_{0}\in\Omega$. Take $\varphi=\varphi(t,x)\in\mathscr{C}_{0}^{\infty}((0,\varepsilon)\times\Omega)$, supported near $(T,x_{0})$. In the operators space $\mathcal{A}^{0}$ consider the local identity partition
@@
Clearly, $\operatorname{Supp}(q_{2})$ is contained in the elliptic set of $T^{*}\mathcal{L}$, so $q_{2}(t,x;D_{t},D_{x})u\in H^{1}(\mathcal{L}_{\varepsilon})$, and $\psi(x,D_{x})q_{2}u\in H^{1}(\mathcal{L}_{\varepsilon})$ .
Let us now examine the first term $q_{1}(t,x;D_{t},D_{x})u$. Here we notice that the composition $\psi(x,D_{x})q_{1}(t,x;D_{t},D_{x})$ provides a well defined global pseudodifferential operator, according to [7, Th. 18.1.35]. In addition, if $\rho=(t,x;\tau,\xi)\in\operatorname{Supp}(\sigma(\psi q_{1}))$, where $\sigma$ denotes the symbol of the operator $\psi q_{1}$, $\rho$ is either an elliptic or hyperbolic point lying in the set $E_{\varepsilon}$, which does not intersect $WF^{1}u$. Therefore, $\psi(x,D_{x})q_{1}(t,x;D_{t},D_{x})u\in H^{1}(\mathcal{L}_{\varepsilon})$.
@@
Here we shall work in a system of local geodesic coordinates $(t,x^{\prime},x_{n};\tau,\xi^{\prime},\xi_{n})$ with $\partial\Omega=\{x_{n}=0\}$ and $\Omega=\{x_{n}>0\}$, see Section 3.1. Hence we will set $\rho_{0}=(T,x^{\prime}_{0},\tau_{0},\xi^{\prime}_{0})$. Recall that the operators of $\mathcal{A}^{0}$ (resp. of $\mathcal{B}{0}$) take the form $q(x_{n},t,x^{\prime},D_{t},D_{x^{\prime}})$ (resp. $\psi(x_{n},x^{\prime},D_{x^{\prime}})$).
As in Case 1 above, we consider $\varphi=\varphi(t,x)\in\mathscr{C}_{0}^{\infty}((0,\varepsilon)\times{\mathbb{R}}^{n})$, supported near $(T,x_{0})$, and a local partition of the identity with tangential pseudodifferential operators, of the form
@@
With notations of Section 3.2, $\operatorname{Supp}(q_{2})\subset\mathcal{H}$, the hyperbolic subset of $T^{*}\partial\mathcal{L}$. Therefore since $\operatorname{Supp}(\psi)\subset\mathscr{R}_{0}(\omega_{T})$, we get $q_{2}(x_{n},t,x^{\prime};D_{t},D_{x^{\prime}})\psi(x_{n},x^{\prime},D_{x^{\prime}})u\in H^{1}(\mathcal{L}_{T})$ .
Finally, as in Case 1, we notice that the composition $q_{1}(x_{n},t,x^{\prime};D_{t},D_{x^{\prime}})\psi(x_{n},x^{\prime},D_{x^{\prime}})$ provides a well defined global tangential pseudodifferential operator, see [7, Th. 18.1.35], whose support is contained in $\mathcal{E}\cup\mathscr{R}(\omega_{T})$ . And this yields $q_{1}(x_{n},t,x^{\prime};D_{t},D_{x^{\prime}})\psi(x_{n},x^{\prime},D_{x^{\prime}})u\in H^{1}(\mathcal{L}_{T})$ according to item (1).
Let us now examine the regularity of the trace $(u_{0},u_{1})$. For this, consider a function $h(t)\in\mathscr{C}^{\infty}_{0}({\mathbb{R}})$, $h(t)=1$ for $|t|\leqslant\varepsilon/2$ and $h(t)=0$ for $|t|\geqslant 3/4\varepsilon$. For $u(t,x)$ solution to (1.1), the function $v=h(t)\psi(x,D_{x})u$ satisfies the wave system
@@
The right hand side of this equation lies in $L^{2}((0,\varepsilon)\times\Omega)$ according to the argument above. Therefore $(v(0,\cdot),\partial_{t}v(0,\cdot))=\psi(x,D_{x})(u_{0},u_{1})$ belongs to $H^{1}_{0}(\Omega)\times L^{2}(\Omega)$. This ends the proof of item (2) in Lemma 3.7, since $\psi(x,D)$ is any operator in $\mathcal{B}^{0}$ supported in $\mathscr{R}_{0}(\omega_{T})$. ∎